find three consecutive integers such that the sum of twice the second and three times the third is five less than six times the first
Answer by vicgonzerx(31)

X=FIRST NUMBER
X+1=2ND NUMBER
X+2=3RD NUMBER
2(X+1)+3(X+2)=6X-5
2X+2+3X+6=6X-5
-X=-5-2-6
X=13
X+1=14
X+2=15
13,14,15------ANSWER